OpenStudy (anonymous):

What is the solution for this inequality? -4x < 28

OpenStudy (anonymous):

Divide -4 to 28.

OpenStudy (anonymous):

\[-4x <28\] \[\frac{ -4 }{ -4 }=\frac{ -4 }{ 28 }\]

OpenStudy (anonymous):

What do you get @Wonka112?

OpenStudy (anonymous):

Or you could do \[\frac{ 28 }{ -4 }\]

OpenStudy (anonymous):

a.x < -4 b.x < 4 c.x > -4 d.x > 4

OpenStudy (anonymous):

@vzfreakz

OpenStudy (anonymous):

That's your answer choices?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Hmm. I don't get why my answer isn't one of them.

OpenStudy (anonymous):

Do you have any idea what it could be?

OpenStudy (anonymous):

I'm not quite sure. But try to bump this question again. Probably, some one else could help.

OpenStudy (anonymous):

Alright. Thanks for trying though super appreciated!

OpenStudy (anonymous):

No problem.